2. A solid sample of 7.50 g K2S is dissolved in 500.0 g H2O in a coffee-cup calorimeter. The temperature of the water rises from 25.0 to 28.7 °C. Calculate dH for the dissolution of potassium sulfide in water. Assume the specific heat of the resulting solution is the same as water; 1.00 calories/gram degree C.
Answer #1

Mass of solution = (500.0 +7.50) = 507.50 8. Heat = mx S x dT .. 507.50g x 1.00 cal/g-9 x (28.7 - 25.0) & = 1877.75 cal Mole